Operations Report – April 4th

Crew 227 Operations Report 04-04-2022

SOL: 8

Name of person filing report: Julien Meert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: Not use

Opportunity rover used: Not use

Curiosity rover used: Not use

Perseverance rover used: Not use

General notes and comments: Nothing to report

Summary of Hab operations : Nothing to report


Water (static tank): 312 gallons

Water (loft tank): 27 gallons

Water Meter: 157260,1 units

Static to Loft Pump used : 38 gallons

Static tank pipe heater (on or off): Off

Static tank heater Off

Toilet tank emptied: Yes

Summary of internet: It was not working today but Shannon is moving the Starlink antenna between our habitation, so when you will receive this it will mean that it’s repaired

Summary of suits and radios: Nothing new

Summary of GreenHab operations: Nothing new

WATER USE: 0 gallons

Heater: Off

Supplemental light: Off

Harvest: 0

Summary of ScienceDome operations: We will let the incubator ON during the night for an experimentation

Dual split: Off

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing special, have a good evening !

Eva Report – April 4th

Crew 227 EVA Report 04/04/2022
SOL: 8
Name of person filing report: Jean Jacobs
Author: Jean Jacobs

Hello capcom,

Thank you for reading us!

Purpose of EVA: Experiment of broken leg fixator on EVA and exploration of the area North near to the HAB.
Start time: 9:00
End time: 12:00
Narrative: After walking to Schubert pass, the EVA team set up for Julie’s experiment. 2 broken legs were brought with us. Cyril and Jean were able to fix their respective leg. The team then walked to Hab view North as well as towards the North to take pictures.
Destination: Schubert Pass and HAB Ridge Rd
Coordinates: 4251500 N; 517900 E
Participants: Jean, Cyril, Sirga, Julie, Ignacio
Roads and routes per MDRS map: Path 1103
Mode of travel: by foot

Jean Jacobs

Sol Summary – April 4th

Crew 227 Sol Summary Report 04/04/22
Sol: 8

Hello Capcom,

Thank you for reading us!

Summary Title: First of 3 leg repairing EVA proved to be tricky, good food for the morale.

Author’s name: Jean Jacobs (XO)

Mission Status: Wonderful news coming from the drone images. Good management of food reserves means a comfortable end of tour.

Sol activity summary:
08:00 Breakfast
09:00 EVA with fractured leg repairing session
13:00 Lunch
14:00 Data analysis and Julie’s experiment fixing legs for Audrey and Cyril
16:00 Team gathering around a banana bread
18:00 Report redaction
19:00 Capcom

Anomalies in work: None

Weather: Sunny morning and windy afternoon

Crew Physical Status: Crew is in shape

EVA: First leg repairing EVA proved to be tricky, beautiful panorama on HAB Ridge Rd

Reports to be filed:
– Crew commander
– Operation report
– Astronomer
– GreenHab officer
– Journalist
– Picture of the day
– Some pictures
– EVA request
– EVA summary

Support Requested: None at this time

Jean Jacobs

Copyright © The Mars Society. All rights reserved. | Main Site